OpenStudy (anonymous):

How to find all points P on the parabola y = x^2 such that the tangent line at P passes through the point ( 0, 4 )?? What if the point were changed to ( 2, 5 ) ?

OpenStudy (anonymous):

doesn't look possible to me

OpenStudy (anonymous):

o.O its a homework question...so i hope its possible

OpenStudy (anonymous):

are sure you wrote the problem correctly?

OpenStudy (anonymous):



OpenStudy (anonymous):

eh sorry! (0,-4)

OpenStudy (anonymous):

can't see how to draw a line though (0,4) tangent to the curve. maybe my thinking is off

OpenStudy (anonymous):

then (2,-5) :/ sorry

OpenStudy (anonymous):

ok so you know that the derivative of \[y=x^2\] is \[y'=2x\] and therefore any point on the curve \[(x,x^2)\] that is tangent has to satisfy \[2x=\frac{x^2+4}{x-0}\]

OpenStudy (anonymous):

that is the slopes must match up

OpenStudy (anonymous):

solve this equation for x to get your answer

OpenStudy (anonymous):

\[2x^2=x^2+4\] \[x^2=4\] \[x=\pm2\] so you have two lines tangent to the curve at (0,-4)

OpenStudy (anonymous):

the one through \[(2,4),(0,-4)\] and the one through \[(-2,4) ,(0,-4)\]

OpenStudy (anonymous):

second one is the same. solve \[2x=\frac{x^2+5}{x-2}\]for x

OpenStudy (anonymous):

im a little confused on how you get those equations where 2x=x^2+5/x-2 and the other